# Combining Constrained Diffusion Models and Numerical Solvers for Efficient and Robust Non-Convex Trajectory Optimization

### v3

@@ 4.2 Constrained Diffusion Model
The original diffusion models (DMs) aim to learn the distribution of locally optimal solutions $p(x^{*}|y)$ through the loss function in Eq. (5) without knowing any constraint information. As a result, the sampled solutions could be close to the groundtruth but still largely violate the constraints. To address this issue, we propose a novel constrained diffusion model (CDM) to minimize the constraint violations in the sampled trajectories. Inspired by the equation loss in PINN [37], we define a violation function $V(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for differentiable constraints in Eq. (1):
@@
where $V(\cdot,\cdot)\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{k};\mathbb{R})$ (differentiable except at $g_{i}=0$ or $h_{j}=0$) maps from a sample $x$ and parameter $y$ and outputs the total constraint violation as a scalar value. In fact, each constraint term can be further customized with different weights that allow different treatments. The violation value is expected to decrease during the training process of diffusion models. We introduce the newly proposed loss function and the corresponding training process as follows.
CDM Training.
We define the locally optimal solution $x^{*}=x_{0}$ as initial sample in the diffusion process with diffusion step $k=0$. By reformulating the forward sampling process as in Eq. (3), we have $x_{0}=\frac{x_{k}-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon}{\sqrt{\widebar{\alpha}_{k}}},\varepsilon\sim\mathcal{N}(0,\mathbf{I})$. The noise $\varepsilon$ is approximated as $\varepsilon_{\theta}(x_{k},k,y)$ with neural-network parameterization. Therefore, we can predict initial sample as $\widehat{x}_{0}=\frac{x_{k}(x_{0},\varepsilon)-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon_{\theta}(x_{k},k,y)}{\sqrt{\widebar{\alpha}_{k}}}$. Then we introduce a constraint violation loss $\mathcal{L}_{\text{vio}}$ on the predicted $\widehat{x}_{0}$ from $x_{k}$, $\forall k\in[K]$:
@@
where $x_{0}^{\text{min}}$ and $x_{0}^{\text{max}}$ are the lower and upper bound for $x_{0}\in\mathcal{X}$. The reason for the clipping is that at the early phase of the training, the parameterized $\varepsilon_{\theta}$ as an approximation of $\varepsilon$ is not accurate thus the predicted $\widehat{x}_{0}$ may not lie on the original range of the data $x_{0}$. Therefore, we clip $\widehat{x}_{0}$ to maintain its range, which guarantees the valid calculation of violation function $V(\cdot,\cdot)$. The scheduling factor $\frac{1}{k}$ regulates the scale of the violation function $V$ for different $x_{k},k\in[K]$. The reason is that the predicted $\widehat{x}_{0}$ will be more noisy when $k$ is large therefore the estimation of violation function can be less reliable. This may not be a theoretically principled choice but is found effective and convenient in our experiments.
@@
Finally, our proposed constrained diffusion training loss is a combination of the original diffusion model training loss and the above violation loss:
@@
Then we plug $\widehat{\varepsilon}_{\theta}$ into $\varepsilon_{\theta}$ in the sampling process [21, 22], and through iterative sampling $x_{k-1}$ from $x_{k}(k=K,\dots,1)$ with initial random noise $x_{K}\sim\mathcal{N}(0,\mathbf{I})$, it produces the generated samples $x_{0}$. This diffusion samples serve as the predicted initial guess in Sec. 4.1 for a new task: $\tilde{x}^{*}_{\text{new}}:=x_{0}$.
@@ 5 Experiment
In this section, we evaluate the proposed framework combining diffusion models and numerical solvers on three trajectory optimization tasks, and compare with several baselines. Our methods include two DMs: the vanilla diffusion model (DM) and the constrained diffusion model (CDM) as proposed in Sec. 4.2. The experiments involve three stages: 1. Collect solutions provided by numerical solvers on each task domain; 2. For data-driven methods, train the models to fit the solution dataset; 3. For unseen new tasks within each domain, sample initial solutions from models and feed into numerical solvers for deriving final solutions. The details of data collection and model training refer to Appendix A.
Tasks.
Our experiments are conducted on three trajectory optimization tasks across different domains with examples, as shown in Fig. 1. With an increasing complexity (non-convexity) of the problem, we introduce: (i). tabletop manipulation for object relocation with a robotic arm (Sec. 5.1); (ii). the two-car reach-avoid problem as a two-player motion planning game (Sec. 5.2); and (iii) the cislunar transfer problem for optimal control of the spacecraft (Sec. 5.3). Each task has various DoF, which is derived with the number of control variables at each time step multiplied with the number of discretized steps along the trajectory. The tabletop manipulation has a higher DoF due to a larger number of time stesp ($T=80$), however, the task is less complex as the cislunar transfer, which has more complicated dynamics thus higher non-convexity. More details for each task refer to Appendix A.
@@
We assume fully known dynamics and environment for the open-loop optimal control problem. To solve the tabletop manipulation and two-car reach-avoid problem, we formulate and solve the problem with Sparse Nonlinear OPTimizer (SNOPT) [19] as the solver $\pi$, which uses the SQP method and supports warm-starting. For the cislunar transfer, we formulate and solve the problem with pydylan, a Python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) [3] together with SNOPT. The detailed problem formulation and solving for each task are included in Appendix A.
@@
For comparison with the proposed framework, we introduce the following baselines:
  • Uniform: This method solely uses a numerical solver with uniformly sampled initial guesses $x\in\mathcal{X}$ for the non-convex trajectory optimization problem. It solves problems from scratch without leveraging any problem-specific prior information.
  • Optimal Uniform: This method uniformly samples from collected locally optimal solutions $\{x^{*}\}$ as initial guesses for the numerical solver, without using any machine learning method to approximate the solution distribution.
  • CVAE_LSTM: This method [32] combines the convolutional variational auto-encoder (CVAE) with long short-term memory (LSTM) to learn the conditional distribution of locally optimal solutions. Specifically, a CVAE with Gaussian Mixture Model (GMM) prior is first used to sample the non-control variables, e.g. time or mass variables, and then an LSTM is adapted to sample the control sequence.
Although there exist other graph-based path planning methods like Rapidly-exploring Random Trees (RRT) [31], A [20], etc., they can not account for both complex dynamical and environment constraints and flexible optimization objectives in the optimal control problem. Consequently, the comparison against them is not provided in our experiments.
@@
Apart from the time consumption for deriving feasible solutions, we also apply diversity measure for each method since there exist multiple local optima for each task. Specifically, the Vendi Score [18] is applied to measure the diversity of the model outputs. The similarity kernel we choose is $k(x,y)=e^{-||x-y||}$.
5.1 Tabletop Manipulation
Task Setting.
For a robotic arm manipulating objects on the tabletop as shown in Fig. 7 in the Appendix, it is important to plan a desired trajectory with goal-reaching and obstacle avoidance. We define the states of the system to be the location of the gripper: $(p_{x},p_{y})\in[0,10]^{2}$. In a simplified simulation setting, the movement of the gripper is modeled with linear dynamics $\dot{p}_{x}=v_{x},\dot{p}_{y}=v_{y}$, where $u=(v_{x},v_{y}),v_{x},v_{y}\in[-1,1]$ are two-dimensional speed control variable. In this task $\mathcal{P}_{y}$, we aim to minimize the time for the gripper to reach the goal while avoiding the obstacles.
Figure 3: Raw trajectory prediction output and the solver output for the tabletop manipulation (top row) and two-car reach-avoid problem (bottom row). From left to right, uniform sampling, CVAE_LSTM, DM, CDM and solver refined locally optimal solution from CDM sampled initial guess.
Results.
In the top row of Fig. 3, our proposed constrained diffuse model generates a nearly collision-free trajectory to reach the goal, which is very similar to the final refined solution from the solver. The DM tries to navigate through the cluttered environment but chooses a path that inevitably collides with obstacles. The baseline CVAE_LSTM is not able to account for both goal-reaching and obstacle avoidance. This can be further verified by checking the constraint violation in Table 2, where both DM and CDM achieve significant reduction on constraint violation. CDM has the lowest violation, which justifies the effectiveness of the additional constraint violation loss for training.
When we use our proposed models to warm-start the solver $\pi$, only minor refinement is needed to derive locally optimal trajectories. We compute the ratio of locally optimal solutions obtained from 600 initial guesses and present the corresponding time consumption (involving both sampling time from the model and numerical solver time) statistics in Table 1. As demonstrated in Table 1, the CDM with the solver is 29.6 times faster than the uniform method based on the top $25\%$-quantile of the computational time for deriving the locally optimal solutions. This is further demonstrated in Fig. LABEL:Fig:tabletop_time_histogram with the solving time histogram of locally optimal solutions, where the CDM with the solver starts to obtain many solutions in a shorter solving time. Then Vendi score result in Table 1 verifies the diversity of our model output, where uniform samples intrinsically have the highest diversity. Thus both efficiency and solution quality are achieved with our framework.
5.2 Two-Car Reach-Avoid
Task settings.
The two-car reach-avoid problem is a two-player game where each car tries to find a minimum-time path to reach the goal while avoiding collision with another car and obstacles simultaneously. Specifically, both car1 and car2 follows the same dynamics as follows:
where $(p_{x},p_{y},v,\theta)$ denotes the state of each car. $p_{x},p_{y}$ are the position, $v$ is the speed, and $\theta$ is the orientation of the car. Each car has two-dimensional controls $u^{1}=(a^{1},\omega^{1}),u^{2}=(a^{2},\omega^{2})$, where $a\in[-1,1]$ is the acceleration and $\omega\in[-1,1]$ is the angular speed. The task setup can be seen in Fig. 1.
Table 1: The ratio of locally optimal solutions, the statistics (mean, standard deviation, quantile and median) of total time consumption including both model sampling time and solver time, and Vendi scores for diversity of sampled trajectories before feeding to the solver.
Task Method Ratio Mean($\pm$±STD) 25%-Quantile Median Vendi Score
Tabletop CDM (Ours) 59% 7.29 $\pm$ 16.57 0.50 1.26 5190.70
DM (Ours) 60% 7.51 $\pm$ 17.50 0.73 1.37 5306.18
CVAE_LSTM 53% 8.53 $\pm$ 20.87 1.10 1.55 516.61
Uniform 62% 31.65 $\pm$ 24.55 14.81 22.72 5917.73
Optimal Uniform 77% 12.80 $\pm$ 17.70 4.11 7.17 3708.58
Two-car CDM (Ours) 95% 17.86 $\pm$ 14.28 7.68 13.83 3873.60
DM (Ours) 95% 18.82 $\pm$ 14.33 8.77 15.61 3574.66
CVAE_LSTM 63% 36.24 $\pm$ 20.55 19.45 33.26 1162.72
Uniform 64% 46.17 $\pm$ 20.63 29.62 43.54 5256.75
Optimal Uniform 83% 25.15 $\pm$ 19.75 10.82 20.89 3708.58
Cislunar DM (Ours) 54% 50.81 $\pm$ 70.30 9.97 26.68 5999.96
CVAE_LSTM 30% 141.31 $\pm$ 125.60 38.41 106.48 1403.38
Uniform 18% 199.34 $\pm$ 116.59 114.02 174.35 6000
Optimal Uniform 23% 177.35 $\pm$ 123.22 70.03 153.97 5918.08
Table 2: The constraint violation value for generated samples before feeding to the solver.
Task Method Mean($\pm$±STD)
Tabletop CDM (Ours) 4.60 $\pm$ 5.38
DM (Ours) 4.73 $\pm$ 5.73
CVAE_LSTM 9.11 $\pm$ 8.16
Uniform 32.20 $\pm$ 3.73
Optimal Uniform 65.44 $\pm$ 40.22
Two-car CDM (Ours) 3.00 $\pm$ 15.36
DM (Ours) 10.99 $\pm$ 35.59
CVAE_LSTM 282.72 $\pm$ 186.40
Uniform 545.66 $\pm$ 294.71
Optimal Uniform 3.27 $\pm$ 3.57
Results.
In the bottom row of Fig. 3, the proposed CDM generates the best prediction of the trajectory solution conditioned on obstacle information $y$, which is very close to the final locally optimal solutions refined by the solver. The unconstrained DM samples reasonable trajectory solution prediction, but with much higher constraint violation. The baseline CVAE_LSTM cannot generate solutions toward goal-reaching. Uniformly sampled initial guesses are usually far from a constraint-satisfying trajectory.
We also evaluate the solving time of obtaining those locally optimal solutions $x^{*}$ from the solver $\pi$ for these five methods in Table 1, and visualize the corresponding histogram in Fig. LABEL:Fig:car_solving_time_histogram. Our proposed CDMs obtained the greatest number of locally optimal solutions from 600 samples as initializations. It is also 3.9 times faster than the uniform method according to the $25\%$-quantile of the computational time for deriving the locally optimal solutions.
To further compare the CDM and the unconstrained DM, we also measure the constraint violation (as defined by Eq. (6)) of all the model outputs in Table 2. Our CDM achieves the lowest constraint violation among all data-driven models. Note that the optimal uniform also achieves small violation because in this task the goal conditions are fixed, the pre-collected solutions will only incur obstacle collision costs which are distance to the obstacle boundary and are inherently small.
The Vendi score in Table 1 shows the superiority of using diffusion models to generate diverse outputs compared to the baseline CVAE_LSTM. We also visualize the qualitatively different trajectory solution with the proposed CDM and the corresponding solver in Figure 4.
Figure 4: Visualization of four solutions for one two-car reach-avoid problem, with CDM predicted initial guesses refined by the solver $\pi$.
5.3 Cislunar Transfer
Task settings.
In this task, we consider a minimum-fuel cislunar transfer with the dynamical model to be a Circular Restricted Three-Body Problem (CR3BP) [27]. Assuming the mass of the spacecraft is negligible, CR3BP describes the equation of motions of a spacecraft under the gravitational force from the sun and moon. Let $m_{1}$ be the mass of the Earth and $m_{2}$ be the mass of the moon, and $\mu=m_{2}/(m_{1}+m_{2})$, the dynamics of the spacecraft in the CR3BP is as follows:
where $\widebar{U}(r_{1}(x,y,z),r_{2}(x,y,z))=-\frac{1}{2}\left((1-\mu)r_{1}^{2}+\mu r_{2}^{2}\right)-\frac{1-\mu}{r_{1}}-\frac{\mu}{r_{2}}$ represents the effective gravitational potential, $r_{1}=\sqrt{(x+\mu)^{2}+y^{2}+z^{2})}$ and $r_{2}=\sqrt{(x-(1-\mu))^{2}+y^{2}+z^{2})}$ are the distance from the spacecraft to the sun and moon, respectively. The objective is to minimize fuel expenditure.
Results.
The cislunar transfer is a challenging problem to solve that requires intensive computational time. In Table 1, our method with the unconstrained DM has a huge improvement in the average solving time and the number of locally optimal solutions obtained, compared to the baseline. It is also 7.0 times faster than the uniform method based on the $25\%$-quantile of computational time for deriving the locally optimal solutions. This can be further verified by the solving time histogram in Fig. LABEL:Fig:_cr3bp_solving_time_histogram. This improvement encourages us to set a short cut-off time for the solver and use more time to sample different initial guesses, which further improves the efficiency of obtaining solutions. Finally, the Vendi score shown in Table 1 indicates that our unconstrained DM generates much more diverse predictions than CVAE_LSTM. Since the cislunar transfer task uses a complex numerical integration method for the dynamics constraint which requires extra effort to incorporate it into the neural network, we will leave the CDM for this task as future work.

### v4

@@ 4.2 DiffuSolve+ with Constrained Diffusion Model
The original diffusion models in DiffuSolve aim to learn the distribution of locally optimal solutions $p(x|y),x\in\mathcal{X},y\in\mathcal{Y}$ with $\mathcal{X}$ as the variable space and $\mathcal{Y}$ as the problem parameter space. DiffuSolve is optimized through the loss function as Eq. (5) without knowing any constraint information. As a result, the sampled solutions could be close to the groundtruth but still largely violate the constraints. To address this issue, we propose an improved DiffuSolve+ method with a constrained diffusion model (CDM) to minimize the constraint violations in the sampled trajectories. Inspired by the equation loss in PINN (Raissi et al., 2019), we define a violation function $V(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for differentiable constraints in Eq. (1):
@@
where $V(\cdot,\cdot)\in PC^{1}(\mathbb{R}^{n},\mathbb{R}^{k};\mathbb{R})$ (piecewise differentiable except at $g_{i}=0$ or $h_{j}=0$) maps from a sample $x$ and parameter $y$ and outputs the total constraint violation as a scalar value. In fact, each constraint term can be further customized with different weights that allow different treatments. The violation value is expected to decrease during the training process of diffusion models. We introduce the newly proposed loss function and the corresponding training process as follows.
DiffuSolve+ Training.
For the diffusion process, we try to generate samples from solution set $x_{0}\in\mathcal{X}$. By reformulating the forward sampling process as in Eq. (3), we have $x_{0}=\frac{x_{k}-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon}{\sqrt{\widebar{\alpha}_{k}}},\varepsilon\sim\mathcal{N}(0,\mathbf{I})$. The noise $\varepsilon$ is approximated as the neural-network parameterized model $\varepsilon_{\theta}(x_{k},k,y)$ for condition $y$ at timestep $k$. Replacing it into previous formula, the clean sample at timestep $k=0$ is predicted as $\widehat{x}_{0}=\frac{x_{k}(x_{0},\varepsilon)-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon_{\theta}(x_{k},k,y)}{\sqrt{\widebar{\alpha}_{k}}}$. The benefits of direct $\widehat{x}_{0}$ prediction is that it does not need iterative diffusion sampling thus avoids gradient backpropagation through multi-step sampling. For constrained diffusion generation with DiffuSolve+, we introduce a constraint violation loss $\mathcal{L}_{\text{vio}}$ on the predicted $\widehat{x}_{0}$ from $x_{k}$, $\forall k\in[K]$:
@@
where the clip function preserves the range of $\widehat{x}_{0}$ to be within the lower $x_{0}^{\text{min}}$ and upper bound $x_{0}^{\text{max}}$ for $x_{0}\in\mathcal{X}$, ensuring the valid calculation of violation function $V(\cdot,\cdot)$. The reason for the clipping is that at the early phase of the training, the parameterized $\varepsilon_{\theta}$ as an approximation of $\varepsilon$ is not accurate thus the predicted $\widehat{x}_{0}$ may not lie on the original range of the data $x_{0}$. The scheduling factor $\frac{1}{k}$ regulates the scale of the violation function $V$ for different $x_{k},k\in[K]$, since the predicted $\widehat{x}_{0}$ will be more noisy when $k$ is large and the estimation of violation function can be less reliable. This may not be a theoretically principled choice but is found effective and convenient in our experiments.
@@
Finally, our proposed constrained diffusion training loss is a combination of the original diffusion model training loss in DiffuSolve as Eq. 5 and the above violation loss:
@@
Then we plug $\widehat{\varepsilon}_{\theta}$ into $\varepsilon_{\theta}$ in the sampling process (Ho et al., 2020; Ho & Salimans, 2022), and through iterative sampling $x_{k-1}$ from $x_{k}(k=K,\dots,1)$ with initial random noise $x_{K}\sim\mathcal{N}(0,\mathbf{I})$, it produces the generated samples $x_{0}$. The diffusion sample serves as the predicted initial guess in Sec. 4.1 for a new task: $\tilde{x}^{*}_{\text{new}}:=x_{0}$.
@@ 5 Experiment
In this section, we first evaluate DiffuSolve with a vanilla diffusion model (DM) on three high-dimensional and highly nonlinear trajectory optimization tasks. Secondly, we present the results of DiffuSolve+ with our novel constrained diffusion model (CDM). As a standalone method, the CDM reduces problem constraint violation of the samples compared to DM. Within DiffuSolve+, CDM helps further improve the efficiency of obtaining locally optimal and feasible solutions compared to DiffuSolve.
The experiments mainly involve three stages: 1. Use an NLP solver to collect solutions on each task domain; 2. Train the data-driven models to fit the solution dataset; 3. For unseen new tasks within each domain, sample the initial guesses from each model that warm-start the solver to derive final solutions. The details of data collection and model training refer to Appendix A.
5.1 Task setup
Our experiments are conducted on three complex non-convex trajectory optimization tasks as illustrated in Figure 1. These tasks span diverse domains and present unique challenges: (i). Two-Car Reach-Avoid: Multi-agent game-theoretical problem is difficult to scale and features multiple equilibria. (ii). Quadrotor Navigation: High-dimensional nonlinear dynamical systems lead to computational inefficiency. (iii). Cislunar Transfer: The highly nonlinear chaotic dynamical system makes spacecraft trajectory optimization time-consuming with numerous local optima. The detailed dynamical settings for each task are as follows. The numerical settings can be seen in Appendix A.
Two-Car Reach-Avoid.
This is a two-player game inspired by (Chen et al., 2018). Each car tries to find a minimum-time path to reach the goal while avoiding collision with another car and obstacles, shown in Figure 1. Specifically, both car1 and car2 follows the same dynamics as follows:
where $(p_{x},p_{y},v,\theta)$ denotes the state of each car. $p_{x},p_{y}$ are the position, $v$ is the speed, and $\theta$ is the orientation of the car. Each car has two-dimensional controls $u^{1}=(a^{1},\omega^{1}),u^{2}=(a^{2},\omega^{2})$, where $a\in[-1,1]$ is the acceleration and $\omega\in[-1,1]$ is the angular speed.
Quadrotor Navigation.
This a navigation problem of a 10D quadrotor, inspired by (Herbert et al., 2017). The quadrotor needs to find the minimum time to reach the goal position while avoiding obstacles shown in Figure 1, where the goal and obstacles are randomly sampled. The quadrotor’s dynamics is defined as follows:
$\dot{x}$ $=v_{x},$ $\dot{v}_{x}$ $=g\tan\theta_{x},$ $\dot{\theta}_{x}$ $=-d_{1}\theta_{x}+\omega_{x},$ $\dot{\omega}_{x}$ $=-d_{0}\theta_{x}+n_{0}a_{x},$ $\dot{z}$ $=v_{z},$
$\dot{y}$ $=v_{y},$ $\dot{v}_{y}$ $=g\tan\theta_{y},$ $\dot{\theta}_{y}$ $=-d_{1}\theta_{y}+\omega_{y},$ $\dot{\omega}_{y}$ $=-d_{0}\theta_{y}+n_{0}a_{y},$ $\dot{v}_{z}$ $=K_{T}a_{z}-g$ (10)
where $x,y,z$ are the position of the quadrotor. $\theta_{x}$, $\theta_{y}$ are the pitch and roll angle, and $\omega_{x}$, $\omega_{y}$ are the corresponding rates. $v_{x},v_{y},v_{z}$ are the speed of the quadrotor. $d_{0},d_{1},K_{T},n_{0}$ are system parameters, and $g=9.81$. The controls are $u=(a_{x},a_{y},a_{z})$.
Cislunar Transfer.
In this task, we consider a minimum-fuel cislunar low-thrust transfer with the dynamical model to be a Circular Restricted Three-Body Problem (CR3BP) (Koon et al., 2000). As shown in Figure 1, the example trajectory (orange) starts from the end of a geostationary transfer spiral (blue) and ends at an arc of the invariant manifold (black). Assuming the mass of the spacecraft is negligible, the CR3BP describes the equation of motion of a spacecraft under the gravitational force from the Earth and moon. Let $m_{1}$ be the mass of the Earth and $m_{2}$ be the mass of the moon, and $\mu=m_{2}/(m_{1}+m_{2})$, the dynamics of the spacecraft in the CR3BP is as follows:
where $\widebar{U}(r_{1}(x,y,z),r_{2}(x,y,z))=-\frac{1}{2}\left((1-\mu)r_{1}^{2}+\mu r_{2}^{2}\right)-\frac{1-\mu}{r_{1}}-\frac{\mu}{r_{2}}$ is the effective gravitational potential, $r_{1}=\sqrt{(x+\mu)^{2}+y^{2}+z^{2})}$ and $r_{2}=\sqrt{(x-(1-\mu))^{2}+y^{2}+z^{2})}$ are the distance from the spacecraft to the sun and moon, respectively. This is a chaotic dynamical system.
5.2 Experiment Setup
@@
We assume fully known dynamics and environment for the open-loop trajectory optimization problem. For all three tasks, we formulate and solve the optimization problem with Sparse Nonlinear OPTimizer (SNOPT) (Gill et al., 2005) as the solver $\pi$, which uses the SQP method and supports warm-starting. For the cislunar transfer, we adopt pydylan, a Python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) (Beeson et al., 2022) to formulate the CR3BP dynamics in addition to SNOPT. The detailed problem formulation and solving for each task are included in Appendix A.
@@
For comparison with the proposed DiffuSolve, we introduce the following baselines:
  • Uniform: This method solely uses an NLP solver with uniformly sampled initial guesses $x\in\mathcal{X}$ for the non-convex trajectory optimization problem. It solves problems from scratch without leveraging any problem-specific prior information.
  • Optimal Uniform: This method uniformly samples from collected locally optimal solutions $\{x^{*}\}$ as initial guesses for the NLP solver. It establishes a performance lower bound for data-driven methods since without learning it cannot generalize to new problems.
  • CVAE_LSTM: This method (Li et al., 2023) combines the convolutional variational auto-encoder (CVAE) with long short-term memory (LSTM) to learn the conditional distribution of locally optimal solutions. Specifically, a CVAE with Gaussian Mixture Model (GMM) prior is first used to sample the non-control variables, e.g. time or mass variables, and then an LSTM is adapted to sample the control sequence.
Although there exist graph-based methods for path planning, such as the Rapidly-exploring Random Trees (RRT) (LaValle, 1998), A (Hart et al., 1968), etc., they do not aim to produce optimal solutions, nor handle the case of nonlinear dynamical systems in complex environmentally constrained problems well. Thus, the comparison against them is not provided in our experiments.
@@
We adopt the Vendi Score (Friedman & Dieng, 2022) as a quantitative metric for measuring the diversity of the model samples. The similarity kernel we choose is $k(x,y)=e^{-||x-y||}$.
5.3 DiffuSolve Results
In this section, we evaluate DiffuSolve’s performance with a vanilla diffusion model (DM) against baseline methods, focusing on solution optimality, feasibility, diversity, and computational efficiency.
5.3.1 Optimality and Feasibility
We generate 600 samples from each method and use them as warm starts for the NLP solver. In Table 1, we present the percentage (ratio) of locally optimal and feasible solutions over 600 samples which are fine-tuned by the NLP solver. Note that before finetuning, this ratio is zero for all methods’ raw samples. The optimality and feasibility are verified by checking the KKT condition and the constraint violation based on the task setup in Sec. 5.1. As shown in Table 1, no method is able to directly generate locally optimal and feasible solutions. When using the samples as warm starts for the NLP solver, our proposed DiffuSolve obtains a greater number of locally optimal and feasible solutions than any non-diffusion methods.
We also visualize the sampled trajectories and the corresponding solver-derived trajectories in Figure 3. For the Uniform and CVAE_LSTM method, the generated trajectories are far from feasible, thus many of them cannot serve as good initial guesses for the NLP solver. DiffuSolve already generates near-feasible trajectories and thus makes NLP solver easier to derive locally optimal and feasible solutions.
Table 1: Percentage (%) (ratio) of locally optimal and feasible solutions over 600 samples.
Task Method
DiffuSolve+ DiffuSolve CVAE_LSTM Uniform Optimal Uniform
Two-car 94.83 94.50 63.17 64.17 83.00
Quadrotor 100.00 99.83 99.67 99.17 98.50
Cislunar 54.00 29.17 17.50 23.33
Note: All values are for ’+ Solver’ (measured on solver outputs with initial guesses from model samples). ’- Solver’ values (measured on raw model samples) were 0% for all cases and are omitted for brevity.
Figure 3: In two-car reach-avoid problem: First row: raw samples (no solver); Second row: corresponding solver derived locally optimal and feasible solutions.
5.3.2 Computational Efficiency
For the locally optimal and feasible solutions obtained in Table 1, Table 2 presents the computational time statistics, involving both sampling time from the model and the NLP solver time. In Table 2, the DiffuSolve with a vanilla DM is 2 $\times$ to 11 $\times$ faster than the uniform method based on the top $25\%$-quantile of the computational time for deriving the locally optimal and feasible solutions. Also in the histogram plots in Figure 5, DiffuSolve obtains more solutions than baseline methods within a short period of time. These results coincide with Figure 3 that since the DiffuSolve generates samples that are close to local optima, it takes a shorter time for the NLP solver to converge.
Thus, it motivates us to set a proper cut-off time for the solver in the future, which can allow more initial guesses to be fed. This could further improve the efficiency of DiffuSolve to derive locally optimal and feasible solutions.
Table 2: Computational time (sampling + solving) statistics and Vendi Score.
Task Method Solving Time Vendi Score
Mean($\pm$±STD) 25%-Quantile Median
Two-car DiffuSolve+ (Ours) 17.86 $\pm$ 14.28 7.68 13.83 3873.60
DiffuSolve (Ours) 18.82 $\pm$ 14.33 8.77 15.61 3574.66
CVAE_LSTM 36.24 $\pm$ 20.55 19.45 33.26 1162.72
Uniform 46.17 $\pm$ 20.63 29.62 43.54 5256.75
Optimal Uniform 25.15 $\pm$ 19.75 10.82 20.89 3708.58
Quadrotor DiffuSolve+ (Ours) 6.65 $\pm$ 3.55 4.47 6.05 5458.23
DiffuSolve (Ours) 7.30 $\pm$ 4.79 4.78 6.62 5428.63
CVAE_LSTM 9.05 $\pm$ 8.30 5.00 6.74 309.18
Uniform 16.87 $\pm$ 13.60 9.28 12.48 5985.10
Optimal Uniform 13.94 $\pm$ 12.08 7.66 10.13 5575.83
Cislunar DiffuSolve (Ours) 50.81 $\pm$ 70.30 9.97 26.68 5999.96
CVAE_LSTM 141.31 $\pm$ 125.60 38.41 106.48 1403.38
Uniform 199.34 $\pm$ 116.59 114.02 174.35 6000.00
Optimal Uniform 177.35 $\pm$ 123.22 70.03 153.97 5918.08
5.3.3 Diversity
In this paper, our proposed DiffuSolve aims to learn the solution distribution of the non-convex problem, instead of predicting a single solution. For example, for the two-car reach-avoid problem, there are multiple equilibrium points that are locally optimal and feasible. In Figure 4, we present 4 qualitatively different solutions with DiffuSolve for each task, showing its ability to generate diverse solutions even within a shorter period of time.
The Vendi Score is also evaluated on the samples from each method in Table 2. While the Uniform method certainly generates the most diverse samples, the DiffuSolve is comparable to Optimal Uniform and outperforms CVAE_LSTM in sample diversity. Especially for the quadrotor navigation, though the solving time for CVAE_LSTM is competitive in Table 2, the Vendi Score shows that its generated solutions are not diverse at all.
Figure 4: Diverse trajectory solutions for two-car reach-avoid, quadrotor navigation and cislunar transfer problem with DiffuSolve method, given the same conditional input.
5.4 DiffuSolve+ Results
In this section, we present the results of DiffuSolve+ on the quadrotor navigation and two-car reach-avoid task. The proposed CDM is evaluated both as a standalone method for reducing constraint violations and within the DiffuSolve+ framework for improving computational efficiency.
Since the current cislunar transfer task originally adopted from (Beeson et al., 2022) uses an adaptive-stepsize integrator for the dynamics constraint, it requires extra effort to incorporate it into the neural network and is beyond our scope. We leave the CDM for this task as future work.
5.4.1 Constraint violation
As a standalone method without the NLP solver, our proposed CDM from Sec. 4.2 has fewer constraint violations compared to the vanilla DM and CVAE_LSTM, displayed in Table 3. For the quadrotor navigation, the CDM is able to directly generate feasible solutions more than any other methods. Also in Figure 3 for the two-car reach-avoid problem, the CDM generates a trajectory that is the closest to the locally optimal and feasible solutions - only car 1 cannot reach the goal but it is already near. It will be easy for the NLP solver to close this gap within DiffuSolve+.
Table 3: Left: The constraint violation of the model samples. Right: The ratio of feasible solutions. Both are measured directly on the model samples without the NLP solver. Here DiffuSolve+: constrained diffusion model, DiffuSolve: vanilla diffusion model
Task Method Constraint Violation (No Solver) Feasible Ratio ‰
Mean($\pm$±STD) 25%-Quantile Median (No solver)
Two-car DiffuSolve+ 3.00 $\pm$ 15.36 0.73 1.44 0
DiffuSolve 10.99 $\pm$ 35.59 1.98 4.52 0
CVAE_LSTM 282.72 $\pm$ 186.40 147.09 292.90 0
Quadrotor DiffuSolve+ 3.06 $\pm$ 7.47 0.10 0.38 81.8
DiffuSolve 3.85 $\pm$ 8.52 0.16 0.63 49.6
CVAE_LSTM 20.19 $\pm$ 22.32 3.36 12.96 16.3
5.4.2 Computational Efficiency
For DiffuSolve+ with the proposed CDM, the computational efficiency is further improved as shown in Table 2. The intuition is clear from Figure 3: when the DiffuSolve+ sample is closer to the local optima, it takes the NLP solver an even shorter time to derive the final solution. We also present the histogram of the computational time in Figure 5. Here we visualize the density of locally optimal and feasible solutions obtained within a time range, where DiffuSolve+ in red is able to obtain more solutions than DiffuSolve within a short period of time from 600 initial guesses.
(a) Two-car reach-avoid.
(b) Quadrotor Navigation.
(c) Cislunar transfer.
Figure 5: The histogram of computational time (including model sampling time and solver time) for different methods to find locally optimal and feasible solutions.